# C-Pack: Packaged Resources To Advance General Chinese Embedding

### v2

@@ 2.3 Model Class: C-TEM
We provide a comprehensive class of well-trained embedding models for the community. Our models take a BERT-like architecture, where the last layer’s hidden state of the special token [CLS] is trained to work as the embedding. There are three different scales for the models: large (with 326M parameters), base (with 102M parameters), and small (with 24M parameters). The large-scale model achieves the highest general representation performances, leading the current public-available models by a considerable margin. The small-scale model is also empirically competitive compared with the public-available models and other model options in C-TEM; besides, it is way faster and lighter, making it suitable to handle massive knowledge bases and high-throughput applications. Thanks to the comprehensive coverage of different model sizes, people are presented with the flexibility to trade off running efficiency and representation quality based on their own needs.
@@
As introduced, the models within C-TEM have been well-trained and achieve a strong generality for a wide variety of tasks. Meanwhile, they can also be further fine-tuned if 1) the embeddings are applied for a specific scenario, 2) the training data is presented for the application scenario. It is empirically verified that the fine-tuned model may bring forth a much better performance for its application, compared with its original model in C-TEM, and the fine-tuned models from other general pre-trained encoders, like BERT. In other words, C-TEM not only presents people with direct usage embeddings but also works as a foundation where people may develop more powerful embeddings.
2.4 Training Recipe
@@
The training recipe of C-TEM is completely released to the public along with C-Pack (Figure 2). Our training recipe has three main components: 1) pre-training with plain texts, 2) contrastive learning with C-MTP (unlabeled), and 3) multi-task learning with C-MTP (labeled), whose specifications are made as follows.
@@
$\bullet$ Pre-Training. Our model is pre-trained on massive plain texts through a tailored algorithm in order to better support the embedding task. Particularly, we make use of the Wudao corpora Yuan et al. (2021), which is a huge and high-quality dataset for Chinese language model pre-training. We leverage the MAE-style approach presented in RetroMAE Liu and Shao (2022); Xiao et al. (2023), which is simple but highly effective. The polluted text is encoded into its embedding, from which the clean text is recovered on top of a light-weight decoder:
@@
($\mathrm{Enc}$, $\mathrm{Dec}$ are the encoder and decoder, $X$, $\tilde{X}$ indicate the clean and polluted text.)
@@
($p$ and $q$ are the paired texts, $q^{\prime}\in Q^{\prime}$ is a negative sample, $\tau$ is the temperature). One critical factor of contrastive learning is the negative samples. Instead of mining hard negative samples on purpose, we purely rely on in-batch negative samples Karpukhin et al. (2020) and resort to a big batch size (as large as 19,200) to improve the discriminativeness of the embedding.
model
Dim
Retrieval
STS
Pair CLF
CLF
Re-rank
Cluster
Average
Text2Vec (base)
768
38.79
43.41
67.41
62.19
49.45
37.66
48.59
Text2Vec (large)
1024
41.94
44.97
70.86
60.66
49.16
30.02
48.56
Luotuo (large)
1024
44.40
42.79
66.62
61.0
49.25
44.39
50.12
M3E (base)
768
56.91
50.47
63.99
67.52
59.34
47.68
57.79
M3E (large)
1024
54.75
50.42
64.30
68.20
59.66
48.88
57.66
Multi. E5 (base)
768
61.63
46.49
67.07
65.35
54.35
40.68
56.21
Multi. E5 (large)
1024
63.66
48.44
69.89
67.34
56.00
48.23
58.84
OpenAI-Ada-002
1536
52.00
43.35
69.56
64.31
54.28
45.68
53.02
BGE (small)
512
63.07
49.45
70.35
63.64
61.48
45.09
58.28
BGE (base)
768
69.53
54.12
77.50
67.07
64.91
47.63
62.80
BGE (large)
1024
71.53 54.98 78.94 68.32 65.11
48.39
63.96
Table 2: Performance of various models on C-MTEB.
$\bullet$ Task-specific fine-tuning. The embedding model is further fine-tuned with C-MTP (labeled). The labeled datasets are smaller but of higher quality. However, the contained tasks are of different types, whose impacts can be mutually contradicted. In this place, we apply two strategies to mitigate this problem. On one hand, we leverage instruction-based fine-tuning Su et al. (2022); Asai et al. (2022), where the input is differentiated to help the model accommodate different tasks. For each text pair ($p$, $q$), a task specific instruction $I_{t}$ is attached to the query side: $q^{\prime}\leftarrow q+I_{t}$. The instruction is a verbal prompt, which specifies the nature of the task, e.g., “search relevant passages for the query”. On the other hand, the negative sampling is updated: in addition to the in-batch negative samples, one hard negative sample $q^{\prime}$ is mined for each text pair ($p$, $q$). The hard negative sample is mined from the task’s original corpus, following the ANN-style sampling strategy in Xiong et al. (2020).
3 Experiments
We conduct experimental studies for the following purposes. P1. The extensive evaluation of different Chinese text embeddings on C-MTEB. P2. The empirical verification of the text embeddings by C-TEM. P3. The exploration of the practical value brought by C-MTP. P4. The exploration of the impacts introduced by the training recipe.
We consider the following popular Chinese text embedding models as the baselines for our experiments: Text2Vec-Chinese [^4] https://huggingface.co/shibing624 base and large; Luotuo [^5] https://huggingface.co/silk-road/luotuo-bert-medium; M3E [^6] https://huggingface.co/moka-ai base and large; multilingual E5 Wang et al. (2022b) and OpenAI text embedding ada 002777https://platform.openai.com/docs/guides/embeddings. The main metric presented in Section 2.1 is reported for each task in C-MTEB.
3.1 General Evaluation
We extensively evaluate C-TEM against popular Chinese text embeddings on C-MTEB as shown in Table 2.888Our C-TEM models are named BGE in the tables. We make the following observations.
First, our models outperform existing Chinese text embeddings by large margins. There is not only an overwhelming advantage in terms of the average performance, but also notable improvements for the majority of tasks in C-MTEB. The biggest improvements are on the retrieval task followed by STS, pair classification, and re-ranking. Such aspects are the most common functionalities of text embeddings, which are intensively utilized in applications like search engines, open-domain question answering, and the retrieval augmentation of large language models. Although the advantages for classification and clustering tasks are not as obvious, our performances are still on par or slightly better than the other most competitive models. The above observations verify the strong generality of C-TEM. Our models can be directly utilized to support different types of application scenarios.
@@ 3.1 General Evaluation
model
Dim
Retrieval
STS
Pair CLF
CLF
Re-rank
Cluster
Average
M3E (large)
1024
54.75
50.42
64.30
68.20
59.66
48.88
57.66
OpenAI-Ada-002
1536
52.00
43.35
69.56
64.31
54.28
45.68
53.02
w.o. Instruct
1024
70.55
53.00
76.77
68.58
64.91
50.01
63.40
BGE-$i$
1024
63.90
47.71
61.67
68.59
60.12
47.73
59.00
BGE-$i$ w.o. pre-train
1024
62.56
48.06
61.66
67.89
61.25
46.82
58.62
BGE-$f$
1024
71.53 54.98 78.94
68.32
65.11
48.39
63.96
Table 3: Ablation of the training data, C-MTP, and the training recipe.
3.2 Detailed Analysis
We investigate the detailed impact of C-MTP and our training recipe. The corresponding experiment results are presented in Table 3 and Table 4.
First of all, we analyze the impact of our training data, C-MTP. As mentioned, C-MTP consists of two parts. 1) C-MTP (unlabeled), which is used for general-purpose fine-tuning; the model produced from this stage is called the intermediate checkpoint, denoted as BGE-$i$. 2) C-MTP (labeled), where the task-specific fine-tuning is further conducted on top of BGE-$i$; the model produced from this stage is called the final checkpoint, noted as BGE-$f$. Based on our observations from the experimental result, both C-MTP (unlabeled) and C-MTP (labeled) substantially contribute to the embedding’s quality.
Regarding C-MTP (unlabeled), despite mostly being curated from unlabeled corpora, this dataset alone brings forth strong empirical performance for the embedding models trained on it. Compared with other baselines like Text2Vec, M3E, and OpenAI text embedding, BGE-$i$ already achieves a higher average performance. A further look into the performances reveals more details. On one hand, C-MTP (unlabeled) makes a major impact on the embedding’s retrieval quality, where BGE-$i$ notably outperforms the baselines in this attribute. On the other hand, the general capability of embedding is primarily established with C-MTP (unlabeled), as BGE-$i$’s performance is close to the baselines on the rest of the aspects, like STS and Clustering. This puts our embedding models in a very favorable position for further improvements.
As for C-MTP (labeled), the dataset is much smaller but of better quality. With another round of fine-tuning on C-MTP (labeled), the empirical advantage is significantly expanded for the final checkpoint BGE-$f$, where it gives rise to a jump in average performance from 59.0 (BGE-$i$) to 63.96 (BGE-$f$). Knowing that the text pairs in C-MTP (labeled) are mainly gathered from retrieval and NLI tasks, the most notable improvements are achieved on closely related tasks, namely retrieval, re-ranking, STS, and pair classification. On other tasks, it preserves or marginally improves performance. This indicates that a mixture of high-quality and diversified labeled data is able to bring forth substantial and comprehensive improvements for a pre-trained embedding model.
We further explore the impact of our training recipe, particularly contrastive learning, task-specific fine-tuning, and pre-training.
One notable feature of our training recipe is that we adopt a large batch size for contrastive learning. According to previous studies, the learning of the embedding model may benefit from the increasing of negative samples Izacard et al. (2021); Qu et al. (2020); Muennighoff (2022). Given our dependency on in-batch negative samples, the batch size needs to be expanded as much as possible. In our implementation, we use a compound strategy of gradient checkpointing and cross-device embedding sharing Gao et al. (2021b), which results in a maximum batch size of 19,200. By making a parallel comparison between bz: 256, 2028, 19,200, we observe consistent improvement in embedding quality with the expansion of batch size (noted as bz). The most notable improvement is achieved in retrieval performance. This is likely due to the fact that retrieval is usually performed over a large database, where embeddings need to be highly discriminative.
Another feature is the utilization of instructions during task-specific fine-tuning. The task-specific instruction serves as a hard prompt. It differentiates the embedding model’s activation, which lets the model better accommodate a variety of different tasks. We perform the ablation study by removing this operation, noted as “w.o. Instruct”. Compared with this variation, the original method BGE-f gives rise to better average performance. Besides, there are more significant empirical advantages on retrieval, STS, pair classification, and re-rank. All these perspectives are closely related to the training data at the final stage, i.e. C-MTP (labeled), where the model is fine-tuned on a small group of tasks. This indicates that using instructions may substantially contribute to task-specific fine-tuning.
One more characteristic is that we use a specifically pre-trained text encoder to train C-TEM, rather than using common choices, like BERT and RoBERTa Liu et al. (2019). To explore its impact, we replace the pre-trained text encoder with the widely used Chinese-RoBERTa999huggingface.co/hfl/chinese-roberta-wwm-ext-large, noted as “BGE-$i$ w.o. pre-train”. According to the comparison with BGE-$i$, the pre-trained text encoder notably improves the retrieval capability, while preserving similar performances on other aspects.
Batch Size
256
2,048
19,200
Retrieval
57.25
60.96
63.90
STS
46.16
46.60
47.71
Pair CLF
62.02
61.91
61.67
CLF
65.71
67.42
68.59
Re-rank
58.59
59.98
60.12
Cluster
49.52
49.04
47.73
Average
56.43
57.92
59.00
Table 4: Impact of batch size.
4 Related Work
The importance of general text embedding is widely recognized, not only for its wide usage in typical applications, like web search and question answering Karpukhin et al. (2020) but also due to its fundamental role in augmenting large language models Lewis et al. (2020); Guu et al. (2020); Borgeaud et al. (2022); Izacard et al. (2022); Shi et al. (2023). Compared with the conventional task-specific methods, the general text embedding needs to be extensively applicable in different scenarios. In recent years, there has been a continual effort in this field, where a series of well-known works are proposed, like Contriever Izacard et al. (2021), GTR Ni et al. (2021b), sentence-T5 Ni et al. (2021a), Sentence-Transformer Reimers and Gurevych (2019), E5 Wang et al. (2022a), OpenAI text embedding Neelakantan et al. (2022), etc. Although it remains an open problem, recent studies highlight the following important factors. Firstly, the training data is desired to be large-scale and diversified, from which the embedding model can learn to recognize different kinds of semantic relationships Izacard et al. (2021); Wang et al. (2022b); Neelakantan et al. (2022). Secondly, the embedding model must be scaled up, as large text encoders are more generalizable across different application scenarios Muennighoff (2022); Ni et al. (2021b, a) in line with observations for the importance of scaling LLMs (Hoffmann et al., 2022; Rae et al., 2021; Brown et al., 2020; Chowdhery et al., 2022; Srivastava et al., 2022; Gao et al., 2021a; Li et al., 2023a; Allal et al., 2023; Muennighoff et al., 2023b). Thirdly, the training recipe must be optimized through pre-training Liu and Shao (2022); Wang et al. (2022a), negative sampling Izacard et al. (2021); Wang et al. (2022a), and multi-task fine-tuning Su et al. (2022); Asai et al. (2022); Sanh et al. (2021); Wei et al. (2021); Muennighoff et al. (2022b, 2023a); Chung et al. (2022). Aside from the above, it is also critical to establish proper benchmarks to evaluate the generality of text embeddings. Unlike previous task-specific evaluations, like MSMARCO Nguyen et al. (2016), SentEval Conneau and Kiela (2018), it is needed to substantially augment the benchmarks so as to evaluate the embedding’s performance for a wide variety of tasks. One representative work is made by BEIR Thakur et al. (2021); Kamalloo et al. (2023), where the embeddings can be evaluated across different retrieval tasks. It is later extended by MTEB Muennighoff et al. (2022a), where all major aspects of text embeddings can be comprehensively evaluated.
Given the above analysis, it can be concluded that the general text embedding is highly resource-dependent, which calls for a wide range of elements, such as datasets, models, and benchmarks. Thus, the creation and public release of the corresponding resources is crucially important.
5 Conclusion
We present C-Pack to advance progress towards general Chinese embedding. C-Pack consists of three core resources 1) The benchmark C-MTEB, covering 6 major tasks of embeddings and 35 datasets, making it the most comprehensive benchmark to evaluate the generality of Chinese embeddings. 2) The training data C-MTP, curated from massive unlabeled corpora and high-quality labeled datasets. Its unprecedented scale, diversity, and quality contribute to the superior generality of our embedding models. 3) The models C-TEM, which are empirically competitive. Their different sizes provide people with the flexibility to trade off efficiency and embedding quality. The entire training recipe is also provided along with these resources. The public release of C-Pack facilitates the usage of general Chinese embedding and also paves the way for its future advancement.
Appendix A C-MTEB Datasets
Table 5: Overview of datasets in C-MTEB.
Name URL Description Categ. Test        Samples
Classification
AmazonReviewsClassification Muennighoff et al. (2022a); McAuley and Leskovec (2013)
https://hf.co/datasets/mteb/amazon_reviews_multi
Sentiment of Amazon reviews
s2s
5,000
IFlyTek Xu et al. (2020a)
https://hf.co/datasets/C-MTEB/IFlyTek-classification
Long text classification for App descriptions
s2s
2,600
JDReview (https://hf.co/datasets/kuroneko5943/jd21)
https://hf.co/datasets/C-MTEB/JDReview-classification
iPhone reviews
s2s
533
MassiveIntentClassification Muennighoff et al. (2022a); FitzGerald et al. (2022)
https://hf.co/datasets/mteb/amazon_massive_intent
Amazon Alexa virtual assistant utterances annotated with the associated intent
s2s
16,500
MassiveScenarioClassification Muennighoff et al. (2022a); FitzGerald et al. (2022)
https://hf.co/datasets/mteb/amazon_massive_scenario
Amazon Alexa virtual assistant utterances annotated with the associated scenario
s2s
16,500
MultilingualSentiment McAuley and Leskovec (2013)
https://hf.co/datasets/C-MTEB/MultilingualSentiment-classification
Sentiment of Amazon reviews
s2s
3,000
OnlineShopping (https://github.com/SophonPlus/ChineseNlpCorpus/blob/master/datasets/online_shopping_10_cats/intro.ipynb)
https://hf.co/datasets/C-MTEB/OnlineShopping-classification
Sentiment Analysis of User Reviews on Online Shopping Websites
s2s
1,000
TNews Xu et al. (2020a)
https://hf.co/datasets/C-MTEB/TNews-classification
Short Text Classification for News
s2s
10,000
Waimai (https://github.com/SophonPlus/ChineseNlpCorpus/blob/master/datasets/waimai_10k/intro.ipynb)
https://hf.co/datasets/C-MTEB/waimai-classification
Sentiment Analysis of user reviews on takeaway platforms
s2s
1,000
Clustering
CLSClusteringP2P Li et al. (2022)
https://hf.co/datasets/C-MTEB/CLSClusteringP2P
Clustering of titles + abstract from CLS dataset. Clustering of 13 sets, based on the main category.
p2p
10,000
CLSClusteringS2S Li et al. (2022)
https://hf.co/datasets/C-MTEB/C-MTEB/CLSClusteringS2S
Clustering of titles from CLS dataset. Clustering of 13 sets, based on the main category.
s2s
10,000
ThuNewsClusteringP2P Li et al. (2006); Li and Sun (2007)
https://hf.co/datasets/C-MTEB/ThuNewsClusteringP2P
Clustering of titles + abstract from the THUCNews dataset
p2p
10,000
ThuNewsClusteringS2S Li et al. (2006); Li and Sun (2007)
https://hf.co/datasets/C-MTEB/ThuNewsClusteringS2S
Clustering of titles from the THUCNews dataset
s2s
10,000
Pair Classification
Cmnli Xu et al. (2020a, b); Conneau and Kiela (2018); Williams et al. (2017)
https://hf.co/datasets/C-MTEB/CMNLI
Chinese Multi-Genre NLI
s2s
139,000
Ocnli Hu et al. (2020)
https://hf.co/datasets/C-MTEB/OCNLI
Original Chinese Natural Language Inference dataset
s2s
3,000
Reranking
T2Reranking Xie et al. (2023)
https://hf.co/datasets/C-MTEB/T2Reranking
T2Ranking: A large-scale Chinese Benchmark for Passage Ranking
s2p
24,382
MMarcoRetrieval Bonifacio et al. (2021)
https://hf.co/datasets/C-MTEB/Mmarco-reranking
mMARCO is a multilingual version of the MS MARCO passage ranking dataset
s2p
7,437
CMedQAv1 Zhang et al. (2017)
https://hf.co/datasets/C-MTEB/CMedQAv1-reranking
Chinese community medical question answering
s2p
2,000
CMedQAv2 Zhang et al. (2018b)
https://hf.co/datasets/C-MTEB/C-MTEB/CMedQAv2-reranking
Chinese community medical question answering
s2p
4,000
Retrieval
T2Retrieval Xie et al. (2023)
https://hf.co/datasets/C-MTEB/T2Retrieval
T2Ranking: A large-scale Chinese Benchmark for Passage Ranking
s2p
24,832
MMarcoRetrieval Bonifacio et al. (2021)
https://hf.co/datasets/C-MTEB/MMarcoRetrieval
mMARCO is a multilingual version of the MS MARCO passage ranking dataset
s2p
7,437
DuRetrieval Qiu et al. (2022)
https://hf.co/datasets/C-MTEB/DuRetrieval
A Large-scale Chinese Benchmark for Passage Retrieval from Web Search Engine
s2p
4,000
CovidRetrieval Qiu et al. (2022)
https://hf.co/datasets/C-MTEB/CovidRetrieval
COVID-19 news articles
s2p
949
CmedqaRetrieval Qiu et al. (2022)
https://hf.co/datasets/C-MTEB/CmedqaRetrieval
Online medical consultation text
s2p
3,999
EcomRetrieval Long et al. (2022)
https://hf.co/datasets/C-MTEB/EcomRetrieval
Passage retrieval dataset collected from Alibaba search engine systems in e-commerce domain
s2p
1,000
MedicalRetrieval Long et al. (2022)
https://hf.co/datasets/C-MTEB/MedicalRetrieval
Passage retrieval dataset collected from Alibaba search engine systems in medical domain
s2p
1,000
VideoRetrieval Long et al. (2022)
https://hf.co/datasets/C-MTEB/VideoRetrieval
Passage retrieval dataset collected from Alibaba search engine systems in video domain
s2p
1,000
STS
AFQMC Xu et al. (2020a)
https://hf.co/datasets/C-MTEB/AFQMC
Ant Financial Question Matching Corpus
s2s
3,861
ATEC (https://github.com/IceFlameWorm/NLP_Datasets/tree/master/ATEC)
https://hf.co/datasets/C-MTEB/ATEC
ATEC NLP sentence pair similarity competition
s2s
20,000
BQ Chen et al. (2018)
https://hf.co/datasets/C-MTEB/BQ
Bank Question Semantic Similarity
s2s
10,000
LCQMC Liu et al. (2018)
https://hf.co/datasets/C-MTEB/LCQMC
A large-scale Chinese question matching corpus.
s2s
12,500
PAWSX Yang et al. (2019)
https://hf.co/datasets/C-MTEB/PAWSX
Translated PAWS evaluation pairs
s2s
2,000
QBQTC Xu et al. (2020a)
https://hf.co/datasets/C-MTEB/QBQTC
QQ Browser Query Title Corpus
s2s
5,000
STSB Cer et al. (2017)
https://hf.co/datasets/C-MTEB/STSB
Translate STS-B into Chinese
s2s
1,360
STS-22 Muennighoff et al. (2022a)
https://hf.co/datasets/mteb/sts22-crosslingual-sts
Chinese news
p2p
656
Appendix B C-MTP Composition
We mine large-scale pairs of data from various domains. Table 6 shows the details for each data.
data source
type of text pairs
# of pairs
URL
cmrc2018
(query, context)
9,669
https://huggingface.co/datasets/cmrc2018
dureader
(query, context)
96,486
https://github.com/baidu/DuReader
simclue
[…]
388,779
https://github.com/CLUEbenchmark/SimCLUE
csl
(title, abstract)
394,846
https://arxiv.org/abs/2209.05034
amazon_reviews_multi
(title, body)
157,762
https://huggingface.co/datasets/amazon_reviews_multi
wiki_atomic_edits
(sentence, edited sentence)
1,213,688
https://huggingface.co/datasets/wiki_atomic_edits
mlqa
(question, context)
70,594
https://huggingface.co/datasets/mlqa
xlsum
(title, summary) (title, text)
89,505
https://huggingface.co/datasets/csebuetnlp/xlsum
wudao
(title, passage)
37,318,330
https://data.baai.ac.cn/details/WuDaoCorporaText
Misc
60,260,341
Table 6: Details for each dataset. The Misc data comes from the Internet, including QA, paper, and news data.
Appendix C English Models
Using our recipe, we also train a set of English text embedding models presented in Table 7. At the time of writing, our English BGE models are state-of-the-art on the English MTEB benchmark (Muennighoff et al., 2022a) across its 56 datasets. Our models outperform significantly larger models, such as SGPT Bloom which has 7.1 billion parameters (Muennighoff, 2022; Scao et al., 2022a, b). We advance the prior state-of-the-art by an absolute 1.1 (Li et al., 2023b). Our training recipe is the same as for our Chinese models, except for the usage of English data. We first finetune on unsupervised datasets including datasets like Wikipedia, CC-net, StackExchange, Reddit, S2orc, and datasets from sentence-transformers. [^10] https://huggingface.co/datasets/sentence-transformers/embedding-training-data We then further fine-tune on supervised datasets including NLI (Gao et al., 2021c), FEVER (Thorne et al., 2018), NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), Quora, StackExchange Duplicates and MEDI Su et al. (2022).

### v3

@@ 3.3. Model Class: BGE
Even with the full package of training data, it is still challenging to learn general text embeddings due to the expensive training process. In our work, we provide a comprehensive class of well-trained embedding models for the community. Our models are based on the BERT-like architecture (Devlin et al., 2018), which go through three-stage of training (to be discussed in the next section). There are three available scales: large (with 326M parameters), base (with 102M parameters), and small (with 24M parameters). The large-scale model achieves the highest general representation performances, leading the current public-available models by a considerable margin. The small-scale model is also empirically competitive compared with the public-available models and other model options in BGE; besides, it is way faster and lighter, making it suitable to handle massive knowledge bases and high-throughput applications. Thanks to the comprehensive coverage of different model sizes, people are presented with the flexibility to trade off running efficiency and representation quality based on their own needs.
@@
As introduced, the models within BGE have been well-trained and achieve a strong generality for a wide variety of tasks. Meanwhile, they also establish a strong foundation for further fine-tuning. The fine-tuning recipe is well formulated in our training recipe, where all people’s need is the preparation of fine-tuning data. It is empirically verified that the fine-tuned model may bring forth a much better performance for its application, compared with its original model in BGE, and the fine-tuned models from other general pre-trained encoders, like BERT. In other words, BGE not only presents people with direct usage embeddings but also works as a foundation where people may develop more powerful embeddings.
Table 2. Performance of various models on C-MTEB.
model Dim Retrieval STS Pair CLF CLF Re-rank Cluster Average
Text2Vec (base) 768 38.79 43.41 67.41 62.19 49.45 37.66 48.59
Text2Vec (large) 1024 41.94 44.97 70.86 60.66 49.16 30.02 48.56
Luotuo (large) 1024 44.40 42.79 66.62 61.0 49.25 44.39 50.12
M3E (base) 768 56.91 50.47 63.99 67.52 59.34 47.68 57.79
M3E (large) 1024 54.75 50.42 64.30 68.20 59.66 48.88 57.66
Multi. E5 (base) 768 61.63 46.49 67.07 65.35 54.35 40.68 56.21
Multi. E5 (large) 1024 63.66 48.44 69.89 67.34 56.00 48.23 58.84
OpenAI-Ada-002 1536 52.00 43.35 69.56 64.31 54.28 45.68 53.02
BGE (small) 512 63.07 49.45 70.35 63.64 61.48 45.09 58.28
BGE (base) 768 69.53 54.12 77.50 67.07 64.91 47.63 62.80
BGE (large) 1024 71.53 54.98 78.94 68.32 65.11 48.39 63.96
3.4. Training Recipe
@@
The training recipe of BGE is completely released to the public along with C-Pack (Figure 2). Our training recipe has three main components: 1) pre-training with plain texts, 2) contrastive learning with C-MTP (unlabeled), and 3) multi-task learning with C-MTP (labeled). As introduced, the public release of training recipe will not only help with the reproduction, but also benefit the improvement of BGE with continual training and fine-tuning.
@@
$\bullet$ Pre-Training. Our model is pre-trained on massive plain texts through a tailored algorithm in order to better support the embedding task. Particularly, we make use of the Wudao corpora (Yuan et al., 2021), which is a huge and high-quality dataset for Chinese language model pre-training. We leverage the MAE-style approach presented in RetroMAE (Liu and Shao, 2022; Xiao et al., 2023), which is simple but highly effective. The polluted text is encoded into its embedding, from which the clean text is recovered on top of a light-weight decoder:
@@
($\mathrm{Enc}$, $\mathrm{Dec}$ indicate the encoding and decoding operations, $X$, $\tilde{X}$ indicate the clean and polluted text.)
@@
($p$ and $q$ are the paired texts, $q^{\prime}\in Q^{\prime}$ is a negative sample, $\tau$ is the temperature). One critical factor of contrastive learning is the negative samples. Instead of mining hard negative samples on purpose, we purely rely on in-batch negative samples (Karpukhin et al., 2020) and resort to a big batch size (as large as 19,200) to improve the discriminativeness of the embedding.
$\bullet$ Task-specific fine-tuning. The embedding model is further fine-tuned with C-MTP (labeled). The labeled datasets are smaller but of higher quality. However, the contained tasks are of different types, whose impacts can be mutually contradicted. In this place, we apply two strategies to mitigate this problem. On one hand, we leverage instruction-based fine-tuning (Su et al., 2022; Asai et al., 2022), where the input is differentiated to help the model accommodate different tasks. For each text pair ($p$, $q$), a task specific instruction $I_{t}$ is attached to the query side: $q^{\prime}\leftarrow q+I_{t}$. The instruction is a verbal prompt, which specifies the nature of the task, e.g., “search relevant passages for the query”. On the other hand, the negative sampling is updated: in addition to the in-batch negative samples, one hard negative sample $q^{\prime}$ is mined for each text pair ($p$, $q$). The hard negative sample is mined from the task’s original corpus, following the ANN-style sampling strategy in (Xiong et al., 2020).
4. Experiments
In this section, we conduct experimental studies for the exploration of following problems. P1. The extensive evaluation of different Chinese text embeddings on C-MTEB. P2. The empirical verification of the text embeddings by BGE. P3. The exploration of the practical value brought by C-MTP. P4. The exploration of the impacts introduced by the training recipe. We consider the following popular Chinese text embedding models as the baselines for our experiments: Text2Vec-Chinese [^10] https://huggingface.co/shibing624 base and large; Luotuo [^11] https://huggingface.co/silk-road/luotuo-bert-medium; M3E [^12] https://huggingface.co/moka-ai base and large; multilingual E5 (Wang et al., 2022b) and OpenAI text embedding ada 002131313https://platform.openai.com/docs/guides/embeddings. The main metric presented in Section 3.1 is reported for each evaluation task in C-MTEB.
Table 3. Ablation of the training data, C-MTP, and the training recipe.
model Dim Retrieval STS Pair CLF CLF Re-rank Cluster Average
M3E (large) 1024 54.75 50.42 64.30 68.20 59.66 48.88 57.66
OpenAI-Ada-002 1536 52.00 43.35 69.56 64.31 54.28 45.68 53.02
BGE-pretrain 1024 63.90 47.71 61.67 68.59 60.12 47.73 59.00
BGE w.o. pre-train 1024 62.56 48.06 61.66 67.89 61.25 46.82 58.62
BGE w.o. Instruct 1024 70.55 53.00 76.77 68.58 64.91 50.01 63.40
BGE-finetune 1024 71.53 54.98 78.94 68.32 65.11 48.39 63.96
4.1. General Evaluation
We extensively evaluate BGE against popular Chinese text embeddings on C-MTEB as shown in Table 2.141414Our BGE models are named BGE in the tables., where we can make the following observations.
First, our models outperform existing Chinese text embeddings by large margins. There is not only an overwhelming advantage in terms of the average performance, but also notable improvements for the majority of tasks in C-MTEB. The biggest improvements are on the retrieval task followed by STS, pair classification, and re-ranking. Such aspects are the most common functionalities of text embeddings, which are intensively utilized in applications like search engines, open-domain question answering, and the retrieval augmentation of large language models. Although the advantages for classification and clustering tasks are not as obvious, our performances are still on par or slightly better than the other most competitive models. The above observations verify the strong generality of BGE. Our models can be directly utilized to support different types of application scenarios.
@@ 4.1. General Evaluation
Using the same recipe as the Chinese models, we also train a set of English text embedding models presented in Table 5. Besides, our English data was created and released together with C-MTP. It was the first time that such comprehensive training data was made publicly available. At the time of public release, our English BGE models achieved the state-of-the-art performance on the English MTEB benchmark (Muennighoff et al., 2022a) across its 56 datasets. Although there were many strong competitors in the English community, such as E5 (Wang et al., 2022b), SGPT (Muennighoff, 2022), GTE (Li et al., 2023b), GTR (Ni et al., 2021b), and OpenAI Ada-002 (Neelakantan et al., 2022), we were able to notably advance the prior SOTA by an absolute 1.1 points in total average, which further verify the effectiveness of our data curation and training method.
4.2. Detailed Analysis
We investigate the detailed impact of C-MTP and our training recipe. The corresponding experiment results are presented in Table 3 and Table 4, respectively.
First of all, we analyze the impact of our training data, C-MTP. As mentioned, C-MTP consists of two parts. 1) C-MTP (unlabeled), which is used for general-purpose fine-tuning; the model produced from this stage is called the intermediate checkpoint, denoted as BGE-pretrain. 2) C-MTP (labeled), where the task-specific fine-tuning is further conducted on top of BGE-pretrain; the model produced from this stage is called the final checkpoint, noted as BGE-finetune. Based on our observations from the experimental result, both C-MTP (unlabeled) and C-MTP (labeled) substantially contribute to the embedding’s quality.
Regarding C-MTP (unlabeled), despite mostly being curated from unlabeled corpora, this dataset alone brings forth strong empirical performance for the embedding models trained on it. Compared with other baselines like Text2Vec, M3E, and OpenAI text embedding, BGE-pretrain already achieves a higher average performance. A further look into the performances reveals more details. On one hand, C-MTP (unlabeled) makes a major impact on the embedding’s retrieval quality, where BGE-pretrain notably outperforms the baselines in this attribute. On the other hand, the general capability of embedding is primarily established with C-MTP (unlabeled), as BGE-pretrain’s performance is close to the baselines on the rest of the aspects, like STS and Clustering. This puts our embedding models in a very favorable position for further improvements.
As for C-MTP (labeled), the dataset is much smaller but of better quality. With another round of fine-tuning on C-MTP (labeled), the empirical advantage is significantly expanded for the final checkpoint BGE-finetune, where it gives rise to a jump in average performance from 59.0 (BGE-pretrain) to 63.96 (BGE-finetune). Knowing that the text pairs in C-MTP (labeled) are mainly gathered from retrieval and NLI tasks, the most notable improvements are achieved on closely related tasks, namely retrieval, re-ranking, STS, and pair classification. On other tasks, it preserves or marginally improves performance. This indicates that a mixture of high-quality and diversified labeled data is able to bring forth substantial and comprehensive improvements for a pre-trained embedding model.
Table 4. Impact of batch size.
Task Batch Size 256 2,048 19,200
Retrieval 57.25 60.96 63.90
STS 46.16 46.60 47.71
Pair CLF 62.02 61.91 61.67
CLF 65.71 67.42 68.59
Re-rank 58.59 59.98 60.12
Cluster 49.52 49.04 47.73
Average 56.43 57.92 59.00
Table 5. Performance of English Models on MTEB.